# Out of Length Text Recognition with Sub-String Matching

### v2

@@ Abstract
Scene Text Recognition (STR) methods have demonstrated robust performance in word-level text recognition. However, in real applications the text image is sometimes long due to detected with multiple horizontal words. It triggers the requirement to build long text recognition models from readily available short (i.e., word-level) text datasets, which has been less studied previously. In this paper, we term this task Out of Length (OOL) text recognition. We establish the first Long Text Benchmark (LTB) to facilitate the assessment of different methods in long text recognition. Meanwhile, we propose a novel method called OOL Text Recognition with sub-String Matching (SMTR). SMTR comprises two cross-attention-based modules: one encodes a sub-string containing multiple characters into next and previous queries, and the other employs the queries to attend to the image features, matching the sub-string and simultaneously recognizing its next and previous character. SMTR can recognize text of arbitrary length by iterating the process above. To avoid being trapped in recognizing highly similar sub-strings, we introduce a regularization training to compel SMTR to effectively discover subtle differences between similar sub-strings for precise matching. In addition, we propose an inference augmentation strategy to alleviate confusion caused by identical sub-strings in the same text and improve the overall recognition efficiency. Extensive experimental results reveal that SMTR, even when trained exclusively on short text, outperforms existing methods in public short text benchmarks and exhibits a clear advantage on LTB. Code: https://github.com/Topdu/OpenOCR.
@@ 2 Method
Figure 2: Overview of SMTR. SMTR is a lightweight recognizer consisting of two cross-attention-based modules, i.e., sub-string encoder and sub-string matcher. $S$ and $\tilde{S}$ denote a sub-string and it regularized counterpart. $Y_{s}$ denotes the next-previous character set of the sub-string. [R] denotes a randomly selected character from the character set. $E_{s}$ and $E_{I}$ mean sub-string and image embeddings, respectively.
Fig. 2 illustrates the overall architecture of SMTR. Given a text image $X\in\mathbb{R}^{3\times H\times W}$, the image encoder extracts its image embeddings $E_{I}$. A sub-string (S) is encoded into both next and previous queries ($Q_{n}$ and $Q_{p}$) by the sub-string encoder. $Q_{n}$ and $Q_{p}$ are then fed into the sub-string matcher, which implicitly matches the sub-string in the text image and recognizes its next and previous characters bidirectionally. In addition, we discuss issues raised within this new string-matching-based paradigm and propose our solutions.
@@ 2.1 Image Encoder
To accommodate various aspect ratios of the input $X$, we develop a dedicated image encoder termed FocalSVTR to extract visual features as in LISTER (Cheng et al. 2023). Firstly, the input $X$ is divided into patch embeddings ($\in\mathbb{R}^{C_{0}\times\frac{H}{4}\times\frac{W}{4}}$) by two convolution with stride 2. Then, following SVTR (Du et al. 2022), three stages comprising [6, 6, 6] layers of focal modulation (Awais et al. 2023) are stacked. At the end of the first and second stages, $C_{0}$ is extended with a factor of 2 by a convolution. In particular, at the end of the second stage, the convolution downsamples the height of the patch embeddings to $\frac{H}{8}$. Finally, the output features ($\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{4}}$) are flattened to obtain image embeddings $E_{I}$ $\in\mathbb{R}^{C\times\frac{HW}{32}}$, where $C=4C_{0}$.
@@ 2.2 Sub-String Encoder
Sub-String Encoding. Given a sub-string $S$, we employ a lightweight encoding scheme to obtain its feature embedding, the next and previous queries as follows. First, each character in $S$ is converted into a C-dim embedding via a character embedding layer, and the sub-string embedding $E_{s}\in\mathbb{R}^{l_{s}\times C}$ is obtained by feature concatenation. Then, the next token $T_{n}\in\mathbb{R}^{1\times C}$ is introduced, which is a shared and learnable token. $T_{n}$ acts as a query and performs cross-attention with $E_{s}$, obtaining the next query $Q_{n}\in\mathbb{R}^{1\times C}$. Similarity, the previous token $T_{p}\in\mathbb{R}^{1\times C}$ is introduced and generates the previous query $Q_{p}\in\mathbb{R}^{1\times C}$. The above process can be formulated as:
$Q_{z}=\sigma\left(\frac{(T_{z}W^{q})(E_{s}W^{k})^{t}}{\sqrt{C}}\right)(E_{s}W^{v})+T_{z}$ (1)
where $W^{q},W^{k},W^{v}\in\mathbb{R}^{C\times C}$ are learnable weights, $z$ is $n$ or $p$, and $\sigma$ is the Softmax function.
2.3 Sub-String Matcher
With both image and sub-string embeddings, a direct matching way is to predict the sub-string’s position index through feature interaction. However, this might be less effective to long text due to only trained on short text, as seen in previous methods using absolute positional embeddings (Yue et al. 2020; Yu et al. 2020; Bautista and Atienza 2022).
To circumvent this limitation, we propose a sub-string matcher module that employs adjacent characters $Y_{s}$ instead of the absolute position as the prediction targets. The matcher first matches the sub-string within the text image, and then uses the next and previous queries to predict its left and right characters. It gets rid of the global-level absolute positioning and regardless of the text length, concentrating solely on whether the local information within the image matches the sub-string’s embedding. We use the multi-head attention mechanism (Vaswani et al. 2017) to implement this, where $Q_{n}$ or $Q_{p}$ as query and image embeddings $E_{I}$ as key and value. Assuming there are $h$ attention heads, each with different weight matrices $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{C\times C_{h}}$, $i$ denotes the index of the attention head and $C_{h}$ = $\frac{C}{h}$. The matching process can be expressed as:
$\text{Matcher}(Q,E_{I})=\text{MHead}(Q,E_{I},E_{I})$
@@ 2.3 Sub-String Matcher
$A_{i}=\sigma(\frac{(QW_{i}^{q})(KW_{i}^{k})^{t}}{\sqrt{C_{h}}})\in\mathbb{R}^{1\times\frac{HW}{32}}$ (2)
Taking $Q_{n}$ as an example, the sub-string matcher computes the attention map $A_{i}$ of $Q_{n}$ and $E_{I}$, which focuses on the position of the next character of the sub-string, as shown in the attention maps in Fig. 2. Consequently, the output of the sub-string matcher represents the next character feature ($\text{Matcher}(Q_{n},E_{I})\in\mathbb{R}^{1\times C}$), which then undergoes a Classifier ($\in\mathbb{R}^{C\times(N_{c}+1)}$) to generate prediction $\tilde{Y}_{n}\in\mathbb{R}^{1\times(N_{c}+1)}$, i.e., $\tilde{Y}_{n}=\text{Classifier}(\text{Matcher}(Q_{n},E_{I}))$, where ${+1}$ is for the end symbol [E]. In the same way, the previous character prediction is obtained by $\tilde{Y}_{p}=\text{Classifier}(\text{Matcher}(Q_{p},E_{I}))$.
@@ 2.6 Inference Augmentation
The base inference fails to distinguish repeated sub-strings in one image, such as the SMTR w/o IA result in Fig. 5 and Fig. 12 in Supplementary. Tab. 1 shows the statistics of repeated sub-strings in typical short text datasets and LTB. The high percentage of repeated sub-strings (near 10%) in LTB seriously affects the recognition. Consequently, we propose IA that first slices the long text image into three short sub-images and then employs a split-merge scheme for recognition. This process is depicted in Fig. 5 and summarized in Alg. 2 in Supplementary. Specifically, the text image is sliced into three sub-images whose width is halved, i.e., the left sub-image Img1, the right sub-image Img2 and the central sub-image Img3. Note that both Img1 and Img2 are overlapped with Img3. Then, by taking advantage of the bidirectional recognition property of SMTR, we recognize Img1 and Img2 following the next and previous prediction paths, respectively, i.e., Inference(Img1, $T_{n}$) and Inference(Img2, $T_{p}$). The process obtains the results shown in the top line of Fig. 5 in roughly half of the inference iterations, as both predictions can be carried out in parallel. Since the slice may operate on characters and lead to mis-recognition, we pick out recognized sub-strings also appearing in Img3 from both sides according to attention maps (highlighted by yellow in Fig. 5), i.e., $S_{s}$ and $S_{e}$. In the following, we predict Img3 with Inference(Img3, $T_{n}$, $S_{s}$, $S_{e}$), which gets the central characters. Finally, the full result, e.g., the SMTR w/ IA result in Fig. 5, is obtained by a simple post-processing.
It is seen in Tab. 1 that IA drastically reduces the percentage of repeated sub-strings in LTB from nearly 10% to 0.4%, thus significantly alleviating the repeated confusion. Meanwhile, the devised split-merge recognition scheme robustly recognizes long text images. The effectiveness of IA will be demonstrated in Tab. 3 and Fig. 6.
@@ 2.7 Long Text Benchmark
Existing public benchmarks mainly focus on short text recognition and predominantly consist of text with no more than 25 characters. However, they also have a few instances with more than 25 characters. These instances are usually regarded as noise and discarded in developing traditional short text models. Here, we collect these instances and construct a Long Text Benchmark (LTB). LTB contains nearly 4.8k samples whose length is above 25. The composition of LTB is detailed in Tab. 7 in Supplementary. Since all samples are excluded from the training process, LTB establishes a new benchmark exclusively for evaluating the performance of STR models on long text recognition. To assess the impact of length variation on recognition, we divide LTB into three parts based on text length: [26-35], [36-55], and $\geq$ 56. The results of typical STR methods and SMTR are presented in Fig. 1 and Tab. 3, which will be discussed later.
@@ 3.1 Datasets and Implementation Details
We evaluate SMTR on both English and Chinese datasets. For English, our models are trained on Union14M-L (Jiang et al. 2023), which contains about 4 million real-world text images ($L\leq 25$). Then, the models are tested on both LTB and two short text benchmarks: (1) Common benchmarks, i.e., ICDAR 2013 (IC13) (KaratzasAU et al. 2013), Street View Text (SVT) (Wang, Babenko, and Belongie 2011), IIIT5K-Words (IIIT) (Mishra, Karteek, and Jawahar 2012), ICDAR 2015 (IC15) (Karatzas et al. 2015), Street View Text-Perspective (SVTP) (Phan et al. 2013) and CUTE80 (CUTE) (Anhar et al. 2014). For IC13 and IC15, we use the versions with 857 and 1,811 images, respectively. (2) Union14M benchmark, which includes seven challenging subsets: Curve (CUR), Multi-Oriented (MLO), Artistic (ART), Contextless (CTL), Salient (SAL), Multi-Words (MLW) and General (GEN). For Chinese, we use Chinese text recognition (CTR) dataset (Chen et al. 2021), which contains four subsets: Scene, Web, Document ($Doc$) and Hand-Writing ($HW$). We train the model on the whole training set and use Scene validation subset to determine the best model, which is then assessed on the test subsets.
We use AdamW optimizer (Loshchilov and Hutter 2019) with a weight decay of 0.05 for training. The LR is set to $6.5\times 10^{-4}$ and batchsize is set to 1024. One cycle LR scheduler (I. Loshchilov and Hutter 2017) with 1.5/4.5 epochs linear warm-up is used in all the 20/100 epochs, where a/b means a for English and b for Chinese. Regarding the aspect ratio, all images are resized to a maximum pixel size of 32 $\times$ 128 if the aspect ratio is less than 4, otherwise, it is resized to $H$ = 32 and $W$ up to 384. Word accuracy is used as the evaluation metric. Data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed and the maximum text length is set to 25 during training. The size of the character set $N_{c}$ is set to 96 for English and 6625 (Li et al. 2022) for Chinese. All models are trained with mixed-precision on 4 RTX 3090 GPUs.
@@
Figure 6: Inference speeds of different methods. SMTR means not using inference augmentation. For a fair comparison, the same encoder is employed for all the methods and only time consumption of the decoding stage is considered.
Figure 7: Illustration of the predictions of two long text images from LTB and three short text images from Union14M Benchmarks. In addition, their attention maps are visualized in Supplementary for qualitative analysis.
3.2 Ablation Study
Method Common Benchmarks Union14M-L Benchmarks LTB Params (×$\times 10^{6}$)
@@ 3.2 Ablation Study
Table 4: Results on short text benchmarks and LTB when the models are trained on Union14M-L (2023) datasets. ”-” means the model cannot recognize long text directly. AR-STR is an AR variant capable of recognizing long text. * denotes that the results on short benchmarks is obtained by training the model on Union14M-L using the code they released.
Effectiveness of Inference Augmentation (IA). We evaluated IA from both accuracy and efficiency aspects. As shown in Tab. 3, IA largely improves the accuracy of SMTR on LTB, with weighted (W-Avg) and arithmetic (A-Avg) average increasing by 3.99% and 7.17%, respectively. Notably, accuracy gains from IA become more distinct as the text length increases, where 2.42%, 6.63%, and 12.4% improvements are observed across the three subsets. The results can be explained as the higher likelihood of repeating sub-strings with the text length increases. Regarding the inference speed, as illustrated in Fig. 6, IA enhances SMTR’s efficiency in long text recognition substantially. This improvement is mainly due to IA’s split-merge processing, which reduces the iterations in long text recognition remarkably.
Comparison with aligned methods on LTB. As aforementioned, some STR models, despite being designed for short text recognition, are also capable of recognizing long text, e.g., CTC-based and auto-regressive (AR) decoding ones. We align four of them in Tab. 3 with SMTR for comparison. Specifically, we uniformly adopt FocalSVTR as the image encoder, keeping their decoders unchanged, and then training them using the same hyper-parameters detailed in section 3.1. The results are presented in Tab. 3 and Tab. 4, where FocalSVTR is the encoder combining with the naive CTC decoder, AR-STR is the encoder plus with a standard AR decode composed of two transformer units (Vaswani et al. 2017), and denotes our aligned reproductions.
As seen in Tab. 3, SMTR consistently outperforms the compared models in terms of accuracy, and their gaps are sharply enlarged as the text length increases. When inspecting the arithmetic average for $L>35$, SMTR outperforms FocalSVTR, AR-STR and LISTER by 15.98%, 33.06%, and 14.90%, respectively. The first two long text examples in Fig. 7 show that all four comparing methods suffer from the problem of missing characters. In addition, AR-STR and PARSeq are only able to recognize the front characters due to employing absolute positional encoding and only trained on short text. LISTER encounters the problem of circular recognition due to capturing incorrect neighbor characters. In contrast, SMTR effectively recognizes these instances by sub-string matching. These results convincingly verify the superiority of SMTR in recognizing long text and addressing the OOL challenge. Attention maps of the two instances are provided in Supplementary.

### v3

@@ Abstract
Scene Text Recognition (STR) methods have demonstrated robust performance in word-level text recognition. However, in real applications the text image is sometimes long due to detected with multiple horizontal words. It triggers the requirement to build long text recognition models from readily available short (i.e., word-level) text datasets, which has been less studied previously. In this paper, we term this task Out of Length (OOL) text recognition. We establish the first Long Text Benchmark (LTB) to facilitate the assessment of different methods in long text recognition. Meanwhile, we propose a novel method called OOL Text Recognition with sub-String Matching (SMTR). SMTR comprises two cross-attention-based modules: one encodes a sub-string containing multiple characters into next and previous queries, and the other employs the queries to attend to the image features, matching the sub-string and simultaneously recognizing its next and previous character. SMTR can recognize text of arbitrary length by iterating the process above. To avoid being trapped in recognizing highly similar sub-strings, we introduce a regularization training to compel SMTR to effectively discover subtle differences between similar sub-strings for precise matching. In addition, we propose an inference augmentation strategy to alleviate confusion caused by identical sub-strings in the same text and improve the overall recognition efficiency. Extensive experimental results reveal that SMTR, even when trained exclusively on short text, outperforms existing methods in public short text benchmarks and exhibits a clear advantage on LTB.
Codehttps://github.com/Topdu/OpenOCR
@@ 1 Introduction
2 Related Work
Existing STR methods mainly focus on word-level text recognition. Among them, attention-based (Lee and Osindero 2016) methods are intensively studied for their impressive performance. Some of these methods (Bautista and Atienza 2022; Yue et al. 2020; Wang et al. 2021; Du et al. 2023; Fang et al. 2021; Zhang et al. 2023; Fang et al. 2023; Zhao et al. 2024; Du et al. 2025) use learnable position embeddings to accurately learn context information. However, because learnable position embeddings are typically not scalable, these methods can only recognize short text. In contract, methods (Shi et al. 2019; Li et al. 2019; Sheng, Chen, and Xu 2019; Qiao et al. 2021; Zheng et al. 2024; Zhang et al. 2024a) employ LSTM (Hochreiter and Schmidhuber 1997; Chung et al. 2014) or masked self-attention mechanism with sinusoidal positional encoding (Lee and Osindero 2016) to model context. They have some length extrapolation capability. Nevertheless, these models are primarily trained on short text. Therefore their inference ability is limited when dealing with long text, resulting in a significant decrease in recognition accuracy. Additionally, some methods do not rely on well-designed attention-based decoders, such as CTC-based (Graves et al. 2006) methods (Shi, Bai, and Yao 2017; Hu et al. 2020; Du et al. 2022) and length-insensitive LISTER (Cheng et al. 2023). They do not perform well on short text recognition although exhibiting slightly stronger capability in long text recognition.
@@ 3 Method
Figure 2: Overview of SMTR. SMTR is a lightweight recognizer consisting of two cross-attention-based modules, i.e., sub-string encoder and sub-string matcher. $S$ and $\tilde{S}$ denote a sub-string and it regularized counterpart. $Y_{s}$ denotes the next-previous character set of the sub-string. [R] denotes a randomly selected character from the character set.
Fig. 2 illustrates the overall architecture of SMTR. Given a text image $X\in\mathbb{R}^{3\times H\times W}$, the image encoder extracts its image embeddings $E_{I}$. A sub-string (S) is encoded into both next and previous hidden representation ($Q_{n}$ and $Q_{p}$) by the sub-string encoder. $Q_{n}$ and $Q_{p}$ are then fed into the sub-string matcher, which implicitly matches the sub-string in the text image and recognizes its next and previous characters bidirectionally. In addition, we discuss issues raised within this new string-matching-based paradigm and propose our solutions.
@@ 3.1 Image Encoder
To accommodate various aspect ratios of the input $X$, we develop a dedicated image encoder termed FocalSVTR to extract visual features as in LISTER (Cheng et al. 2023). Firstly, the input $X$ is divided into patch embeddings ($\in\mathbb{R}^{C_{0}\times\frac{H}{4}\times\frac{W}{4}}$) by two convolution with stride 2. Then, following SVTR (Du et al. 2022), three stages comprising [6, 6, 6] layers of focal modulation (Yang et al. 2022) are stacked. At the end of the first and second stages, $C_{0}$ is extended with a factor of 2 by a convolution. In particular, at the end of the second stage, the convolution downsamples the height of the patch embeddings to $\frac{H}{8}$. Finally, the output features ($\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{4}}$) are flattened to obtain image embeddings $E_{I}$ $\in\mathbb{R}^{C\times\frac{HW}{32}}$, where $C=4C_{0}$.
@@ 3.2 Sub-String Encoder
Sub-String Encoding. Given a sub-string $S$, firstly, each character in $S$ is converted into a C-dim embedding via a character embedding layer, and the sub-string embedding $E_{s}\in\mathbb{R}^{l_{s}\times C}$ is obtained by feature concatenation. Then, the next token $T_{n}\in\mathbb{R}^{1\times C}$ is introduced, which is a shared and learnable token. $T_{n}$ acts as a query and performs cross-attention with $E_{s}$, obtaining the next sub-string hidden representation $Q_{n}\in\mathbb{R}^{1\times C}$. Similarity, the previous token $T_{p}\in\mathbb{R}^{1\times C}$ is introduced and generates the previous sub-string hidden representation $Q_{p}\in\mathbb{R}^{1\times C}$. Assuming there are $h$ attention heads, each with different parameters $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{C\times C_{h}}$, $i$ denotes the index of the attention head. The above process can be formulated as:
$Q_{z}=\text{MHead}(T_{z},E_{s},E_{s})+T_{z}$
@@
$A_{i}=\sigma\left((QW_{i}^{q})\left(KW_{i}^{k}\right)^{t}C_{h}^{-0.5}\right)\in\mathbb{R}^{1\times\frac{HW}{32}}$ (1)
where $z$ is $n$ or $p$, $C_{h}$ = $\frac{C}{h}$, and $\sigma$ is the Softmax function.
3.3 Sub-String Matcher
We use one cross-attention (Vaswani et al. 2017) to implement sub-string matching, where the sub-string hidden representation $Q_{n}$ or $Q_{p}$ as query and image embeddings $E_{I}$ as key and value. The process can be expressed as:
$\text{Matcher}(Q_{z},E_{I})=\text{MHead}(Q_{z},E_{I},E_{I})$ (2)
Taking $Q_{n}$ as an example, the sub-string matcher computes the attention map $A_{i}$, which focuses on the position of the next character of the sub-string, as shown in the attention maps in Fig. 2. Consequently, the output of the sub-string matcher represents the next character feature $F_{n}=\text{Matcher}(Q_{n},E_{I})\in\mathbb{R}^{1\times C}$, which then undergoes a Classifier ($\in\mathbb{R}^{C\times(N_{c}+1)}$) to generate prediction $\tilde{Y}_{n}\in\mathbb{R}^{1\times(N_{c}+1)}$, i.e., $\tilde{Y}_{n}=\text{Classifier}(F_{n})$, where ${+1}$ is for the end symbol [E]. In the same way, the previous character prediction is obtained by $\tilde{Y}_{p}=\text{Classifier}(F_{p}),F_{p}=\text{Matcher}(Q_{p},E_{I})$.
@@ 3.6 Inference Augmentation
The base inference fails to distinguish repeated sub-strings in one image, such as the SMTR w/o IA result in Fig. 5. Tab. 1 shows the statistics of repeated sub-strings in typical short text datasets and LTB. The high percentage of repeated sub-strings (near 10%) in LTB seriously affects the recognition. Consequently, we propose IA that first slices the long text image into three short sub-images and then employs a split-merge scheme for recognition. This process is depicted in Fig. 5. Specifically, the text image is sliced into three sub-images whose width is halved, i.e., the left sub-image Img1, the right sub-image Img2 and the central sub-image Img3. Note that both Img1 and Img2 are overlapped with Img3. Then, by taking advantage of the bidirectional recognition property of SMTR, we recognize Img1 and Img2 following the next and previous prediction paths, respectively, i.e., Inference(Img1, $T_{n}$) and Inference(Img2, $T_{p}$). The process obtains the results shown in the top line of Fig. 5 in roughly half of the inference iterations, as both predictions can be carried out in parallel. Since the slice may operate on characters and lead to mis-recognition, we pick out recognized sub-strings also appearing in Img3 from both sides according to attention maps (highlighted by yellow in Fig. 5), i.e., $S_{s}$ and $S_{e}$. In the following, we predict Img3 with Inference(Img3, $T_{n}$, $S_{s}$, $S_{e}$), which gets the central characters. Finally, the full result, e.g., the SMTR w/ IA result in Fig. 5, is obtained by a simple post-processing.
It is seen in Tab. 1 that IA drastically reduces the percentage of repeated sub-strings in LTB from nearly 10% to 0.4%, thus significantly alleviating the repeated confusion. The effectiveness of IA will be demonstrated in Tab. 3.
@@ 3.7 Long Text Benchmark
Existing public datasets (Jung et al. 2011; Nagy, Dicker, and Meyer-Wegener 2012; Veit et al. 2016; Shi et al. 2017; Zhang et al. 2017; Liu et al. 2017; Mathew, Jain, and Jawahar 2017; He et al. 2018; Nayef et al. 2019; Sun et al. 2019; Chng et al. 2019; Zhang et al. 2019; Krylov, Nosov, and Sovrasov 2021; Singh et al. 2021; Long et al. 2022) predominantly consist of short text ($L\leq$ 25). However, they also have a few instances with more than 25 characters. These instances are usually regarded as noise and discarded in developing traditional short text models. We collect these instances and construct a Long Text Benchmark (LTB). LTB contains nearly 4.8k samples whose length is above 25. Since all samples are excluded from the training process, LTB establishes a new benchmark exclusively for evaluating the performance of STR models on long text recognition. To assess the impact of length variation on recognition, we divide LTB into three parts based on text length: [26, 35], [36, 55], and $\geq$ 56, as shown in Tab. 3.
@@ 4.1 Datasets and Implementation Details
We evaluate SMTR on both English and Chinese datasets. For English, our models are trained on Union14M-L (Jiang et al. 2023), which contains about 3.2 million real-world text images ($L\leq 25$). Then, the models are tested on both LTB and two short text benchmarks: (1) Common benchmarks, i.e., ICDAR 2013 (IC13) (KaratzasAU et al. 2013), Street View Text (SVT) (Wang, Babenko, and Belongie 2011), IIIT5K-Words (IIIT) (Mishra, Karteek, and Jawahar 2012), ICDAR 2015 (IC15) (Karatzas et al. 2015), Street View Text-Perspective (SVTP) (Phan et al. 2013) and CUTE80 (CUTE) (Anhar et al. 2014). For IC13 and IC15, we use the versions with 857 and 1,811 images, respectively. (2) Union14M-Benchmark, which includes seven challenging subsets: Curve (CUR), Multi-Oriented (MLO), Artistic (ART), Contextless (CTL), Salient (SAL), Multi-Words (MLW) and General (GEN). Note that the 6,600 samples overlapping between Union14M-L and Union14M-Benchmark are filtered (Du et al. 2024). For Chinese, we use Chinese text recognition (CTR) dataset (Chen et al. 2021), which contains four subsets: Scene, Web, Document (Doc) and Hand-Writing (HW). We train the model on the whole training set and use Scene validation subset to determine the best model, which is assessed on the test subsets.
We use AdamW optimizer (Loshchilov and Hutter 2019) with a weight decay of 0.05 for training. The LR is set to $6.5\times 10^{-4}$ and batchsize is set to 1024. One cycle LR scheduler (I. Loshchilov and Hutter 2017) with 1.5/4.5 epochs linear warm-up is used in all the 20/100 epochs, where a/b means a for English and b for Chinese. Regarding the aspect ratio, all images are resized to a maximum pixel size of 32 $\times$ 128 if the aspect ratio is less than 4, otherwise, it is resized to $H$ = 32 and $W$ up to 384. Word accuracy is used as the evaluation metric. Data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed and the maximum text length is set to 25 during training. The size of the character set $N_{c}$ is set to 96 for English and 6625 (Li et al. 2022) for Chinese. The number of attention heads $h$ in sub-string Encoder and Matcher are set to $\frac{C}{32}$ and 2, respectively. All models are trained with mixed-precision on 4 RTX 3090 GPUs.
4.2 Ablation Study
Effectiveness of Inference Augmentation (IA). As shown in Tab. 3, IA largely improves the accuracy of SMTR on LTB, with weighted (W-Avg) and arithmetic (A-Avg) average increasing by 3.99% and 7.17%, respectively. Notably, accuracy gains from IA become more distinct as the text length increases, where 2.42%, 6.63%, and 12.4% improvements are observed across the three subsets. This trend arises because longer texts are more likely to contain repeating sub-strings. IA effectively eliminates these repetitions, as shown in Tab. 1, leading to substantial improvements in recognition accuracy.
@@ 4.2 Ablation Study
Figure 6: Illustration of the predictions of two long text images from LTB and three short text images from Union14M-Benchmark. In addition, their attention maps are visualized in Supplementary for qualitative analysis.
Fair comparison with aligned methods on LTB. As aforementioned, some STR models, despite being designed for short text recognition, are also capable of recognizing long text, e.g., CTC-based and auto-regressive (AR) decoding ones. We align four of them in Tab. 3 with SMTR for comparison. Specifically, we uniformly adopt FocalSVTR as the image encoder, keeping their decoders unchanged, and then training them using the same hyper-parameters detailed in section 3.1. The results are presented in Tab. 3 and Tab. 4, where FocalSVTR is the encoder combining with the naive CTC decoder, AR-STR is the encoder plus with a standard AR decode composed of two transformer units (Vaswani et al. 2017), and denotes our aligned reproductions.
As seen in Tab. 3, SMTR consistently outperforms the compared models in terms of accuracy, and their gaps are sharply enlarged as the text length increases. When inspecting the arithmetic average for $L>35$, SMTR outperforms FocalSVTR, AR-STR and LISTER by 15.98%, 33.06%, and 14.90%, respectively. The first two long text examples in Fig. 6 show that all four comparing methods suffer from the problem of missing characters. In addition, AR-STR and PARSeq are only able to recognize the front characters due to employing absolute positional encoding and only trained on short text. LISTER encounters the problem of circular recognition due to capturing incorrect neighbor characters. In contrast, SMTR effectively recognizes these instances by sub-string matching. These results convincingly verify the superiority of SMTR in recognizing long text and addressing the OOL challenge.
Method Common Benchmarks Union14M-Benchmark LTB Params (×$\times 10^{6}$)
@@
Table 4: Results on short text benchmarks and LTB when the models are trained on Union14M-L (2023) datasets. ”-” means the model cannot recognize long text directly. For short text benchmarks, we use the next inference. * denotes that the results on short benchmarks is obtained by training the model on Union14M-L using the code they released.